1-Company XYZ plans to acquire a new automated welding system to replace the existing manual system. This new system will initially cost $ 600,000 and will be amortized at an ACC rate of 30%. The expected life of the system is 4 years, and the company estimates it will be worth $ 100,000 at the end of this period. Since the new automated system will be more efficient than the old one, the company can expect to achieve a cost savings of $ 180,000 per year before tax during the period. If the ERR is 15% and the tax rate is 44%, what is the net present value of the new system?
1
Expert's answer
2020-03-24T09:35:29-0400

"NPV = -600,000 + \frac{180,000×0.56 + 0.3×600,000}{1.15} + \frac{180,000×0.56 + 0.3×420,000}{1.15^2} + \frac{180,000×0.56 + 0.3×294,000}{1.15^3} + \frac{180,000×0.56 + 0.3×205,800 + 100,000}{1.15^4} =90,045.96."